OpenStudy (anonymous):

@ganeshie8 A toy company took a sample of its toys and found that 3% of the toys were defective. If the company shipped 1200 toys, how many of the toys would the company expect to be defective?

OpenStudy (ilovepie):

1200*.03=36. So about 36 toys would be defective.

OpenStudy (paki):

take 3% of 1200... that will be the answer... @ILovePie is rite....

OpenStudy (anonymous):

Thanks wait

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The table shows the results of drawing one marble from a bag of colored marbles. Each marble was replaced after the outcome was recorded. If there are a total of 50 marbles in the bag, about how many of the marbles are green? red=6 green=4 orange=7 black=3 A. 20 B. 8 C. 10 D. 4

OpenStudy (anonymous):

HELP please

OpenStudy (anonymous):

i think it is 8 @paki

OpenStudy (sleepyjess):

it says how many are green

OpenStudy (anonymous):

oh so 4

OpenStudy (sleepyjess):

unless they want you to figure out how many there would be if they kept the same ratios just 50 marbles not 20

OpenStudy (sleepyjess):

if that is what they want then it would be 10 since 20 *2 1/2 is 50

OpenStudy (sleepyjess):

where is the table?

OpenStudy (paki):

yeah it's 10...

OpenStudy (anonymous):

the table is where it shows red=6

OpenStudy (sleepyjess):

ok

OpenStudy (anonymous):

so ten

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

Ellie has 10 coins in her pocket. She has 3 quarters, 2 pennies, and 5 dimes. If Ellie reaches into her pocket and draw a coin without looking, what is the probability that she will draw a penny? A. 50% B. 25% C. 20% D. 2%

OpenStudy (sleepyjess):

3+2+5=?

OpenStudy (sleepyjess):

10

OpenStudy (anonymous):

10

OpenStudy (sleepyjess):

2/10 are pennies so 2/10=

OpenStudy (anonymous):

2/10=0.2

OpenStudy (sleepyjess):

move the decimal to the right 2 places .2 ---->20

OpenStudy (anonymous):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number less than 5 is rolled? A. 167 B. 833 C. 333 D. 667

OpenStudy (sleepyjess):

4/6 or 2/3 numbers are less than 5 so do 2/3 and tell me what you get

OpenStudy (anonymous):

0.66

OpenStudy (anonymous):

D.

OpenStudy (sleepyjess):

yes